7. A binomial distribution of 8,250 trials has a probability of p 0.638. What is the sta deviation? O 1,905.4 00.2 O 5,263.5 O 43.7 8. The probability of students choosing the left lunch line at school is p 0.431. Find the variance of the binomial distribution of 982 students choosing the left lunch line today 0 423.2 O558.8 015.5 O 240 8 9. A local poll found that only 17% of the 546 people surveyed feel the mayors new city tax is effective. Find the mean of this binomial distribution 77.0 453.2 8.s 92 8
Answer #1

(7)

n = 8250

p = 0.638

q = 1 - p = 0.362

Standard deviation is given by:

vrpi-V8250 × 0.638 × 0.362 43.7

So,

Correct option:

43.7

(8)

n = 982

p = 0.431

q = 1 - p = 0.569

Variance is given by:

s^{2}=npq=982 imes 0.431 imes 0.569=240.8

So,

Correct option:

240.8

(9)

n = 546

p = 0.17

Mean is given by:

-= np-546 × 0.17-92.8

So,

Correct option:

92.8